please help me to solve this problem.
The absolute value of 2x is less than or equal to the absolute of x-3
|2x|<=|x-3|

Answer by tommyt3rd(5050)

We start by finding points from which we will find the boundaries for the intervals

2x=x-3 or 2x=-x+3
x=-3 or x=1

Next we test all intervals

On x<-3 we test x=-4 and see 8<7 (false)

On -3 On x>1 we test x=2 and see 4<1 (false)

Our solution interval: [-3,1]


:)